OpenStudy (anonymous):

f(x) + x^2[f(x)]^3 = 10 f(1)=2 find f'(1)

OpenStudy (anonymous):

please help!

zepdrix (zepdrix):

\(\Large\bf \color{royalblue}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

So derivative huh? Where ya stuck?

OpenStudy (anonymous):

I have f'(x)+2x[f(x)]^3 +x^2 (3[f(x)]^2 * f'(x))

zepdrix (zepdrix):

Ok very good. And the derivative of the `right side` ?

OpenStudy (anonymous):

0

zepdrix (zepdrix):

\[\Large\bf\sf f(x)+x^2\left[f(x)\right]^3\quad=\quad 10\] \[\Large\bf\sf f'(x)+2x\left[f(x)\right]^3+3x^2\left[f(x)\right]^2\;f'(x)\quad=\quad 0\]Ok good, looks like you've got a great so far.

zepdrix (zepdrix):

Now let's evaluate the expression at x=1.

OpenStudy (anonymous):

but f'(1) is unknown

zepdrix (zepdrix):

\[\Large\bf\sf f'(\color{orangered}{1})+2(\color{orangered}{1})\left[f(\color{orangered}{1})\right]^3+3(\color{orangered}{1})^2\left[f(\color{orangered}{1})\right]^2\;f'(\color{orangered}{1})\quad=\quad 0\]

zepdrix (zepdrix):

Correct. After we everything `else` plugged in, we're going to try and `solve for` f'(1). We want to isolate the f'(1).

OpenStudy (anonymous):

how do you isolate f'(1)?

zepdrix (zepdrix):

We'll have to take advantage of factoring and some other simple algebra. Let's not worry about that just yet. Let's get all the coefficients simplified first.

zepdrix (zepdrix):

We need to use the initial data that they gave us,\[\Large\bf\sf f(1)=2\]

OpenStudy (anonymous):

yes I understand that much and after simplifying. f'(1)+16+12*f'(1)=0

OpenStudy (anonymous):

is that correct?

zepdrix (zepdrix):

Mmm good good.

OpenStudy (anonymous):

now I am really stuck

zepdrix (zepdrix):

\[\Large\bf\sf \color{royalblue}{f'(1)}+16+\color{royalblue}{12\;f'(1)}=0\]Combine like-terms.

zepdrix (zepdrix):

If f'(1) is an apple. You currently have one apple + 16 + 12apples = 0

OpenStudy (anonymous):

so 13 apples?

zepdrix (zepdrix):

\[\Large\bf\sf 16+13\;f'(1)=0\]Ok good.

zepdrix (zepdrix):

How do we solve for f'(1) from here? :o

OpenStudy (anonymous):

so 13/16?

zepdrix (zepdrix):

Hmm I think your fraction is upside down +_+ and missing a negative, yes?

zepdrix (zepdrix):

Careful with that quick mental math lol

OpenStudy (anonymous):

lol yeah I should know better than to do mental math

OpenStudy (anonymous):

-1.23?

zepdrix (zepdrix):

f'(1) = -16/13 Yessss good job \c:/

OpenStudy (anonymous):

thank you can you possibly help me with more?? :)

zepdrix (zepdrix):

Close this thread down. Open a new one so we have a nice clean space to work with :O Also, that way in case I get too busy, someone else can see your question at the top of the list :D

OpenStudy (anonymous):

don't help anyone else! Just me please! :)

zepdrix (zepdrix):

Lol there are plenty of smart people who can help you on here XD Don't worry. But if you're waiting long, try typing @zepdrix or @ followed by someone else in the lobby with a 99 by their name. It sends an alert to them so they can quickly get to your question.